# Large earthquakes follow highly unequal ones

### v1

@@ Abstract
It was conjectured for a long time that the tectonic plates are in a self-organized state of criticality and that the Gutenberg-Richter (power) law is a manifestation of that. It was recently shown that for a system near criticality, the inequality of their responses toward external driving could indicate proximity to the critical point. In this work, we show with numerical simulations and seismic data analysis that large earthquake events have a tendency to follow events that are highly unequal. We have applied this framework to various tectonically active regions, such as North America, Southern Japan, parts of South-East Asia and Indonesia.
@@ I Introduction
When tectonic plates move (at a very slow rate), they undergo a stick-slip motion that causes earthquakes when the stored potential energy due to the motion is released due to a sudden slip event. While much is known through more than a century of studies related to the motions and the subsequent energy released, its understanding still faces severe difficulties, particularly due to the inaccessibility of the regions of origin (often tens of kilometers underground) [1]. However, the phenomenon of earthquake has fascinated researchers from various disciplines alike not only due to their devastating consequences, but also due to their remarkable regularities in the statistical sense [2].
@@
Particularly, it is known that the earthquake magnitudes follow the Gutenberg-RIchter law [3], $\log_{10}N=a-bM$, stating that the number of earthquakes ($N$) of magnitude $M$ are related exponentially ($a$ and $b$ are constants). It is also observed that the rate of aftershocks decays with time in a power-law $n(t)=K/(t+c)^{p}$, known as the Omori-Utsu law [4, 5]. There are other empirical laws about earthquake dynamics as well, which are verified using data analysis across various regions in the world. While a complete understanding of the mechanism behind the earthquake phenomenon is still lacking, the empirical observations have been useful in gaining insights into the earthquake dynamics, improving the prospective to model the phenomenon and to also aim for a forecasting possibility, which has so far largely evaded understanding.
@@
There have been some efforts in searching for regularities in earthquake statistics that could potentially lead to hazard analysis. One such area is the change in the exponent value prior to a large event [6, 7, 8]. This methods follows from the stress dependence of avalanche size exponent in sheared rocks, and also in fracture. In such cases, the avalanche size exponent value decreases prior to a large event. In earthquake also, when expressed in terms of energy, the size distributions follow a power law. The values of the power law exponent decreases, if the events are sampled prior to a large event.
@@
In this work, we analyze a regularity, which depends on the inequality of the earthquake energy releases. Particularly, we note that prior to a large earthquake, the energy released are more unequal than when such large events are not happening. Inequality in responses of a driven disordered SOC-like systems is known to be useful in terms of its universal features [9, 10]. It has also shown promises, both in models and (compression) fracture experiments, to have forecasting capabilities of imminent large events [11, 12]. Here we have measured inequality of energy released within a time-window for earthquake time series in various models and also in the data in several regions. It has then been shown to have a positive correlation with ”large” events that immediately follow such a time-window. We have quantified the earthquake inequalities through Gini and Kolkata indices, and then shown its positive correlation with the events that follow.
@@
For the remaining part of the paper, we first describe the models used and the method of measuring the inequalities in their responses. We then correlate the inequalities of the responses with subsequent events. Finally, we compare with data from different regions on earth and show that qualitatively such correlations hold, indicating an underlying principle of relating response inequalities with proximity to a critical point.
@@ II Models and methods
Here we first describe the models used for this study and then move on to define the inequality indices that are used.
II.1 Train model of earthquake
The train model of earthquake is a simplification on the Burridge-Knopoff model [13], where the stick-slip motion of the overlapping tectonic plates are modeled through massive blocks, connected through linear elastic springs, and driven over a rough surface. The frictional force creates the stick-slip motion and the (elastic) energy released can be measured through the motion of the blocks.
We consider a further simplification on the model [14], where we take the frictional surface to be discrete set of blocks, but fixed in space. When the “train” moves over it, slowly pulled by the rightmost block (“engine”), say, some blocks of the train overlaps with the fixed blocks below. At that point, they have a random pinning force that oppose the motion, and is drawn from a uniform distribution in (0,1) randomly. The forces on a block in the train at any given instance are, therefore, the elastic forces from its nearest neighbors due to potential stretching, and the random pinning force. If the pull due to the elastic force exceeds the random pinning, then the block is moved towards the direction of the pulling, hopping over to the next block below, only if that space is not occupied already. The train only moves along the direction of pulling and never opposite.
II.2 Sandpile-like model of earthquake
We adopt the model introduced in ref. [16] for this. Sandpile models typically measure event size by the number of topplings, the present model defines physically meaningful earthquake observables. As a result, a Gutenberg–Richter (Power) law emerges only in a specific parameter regime. We consider a two-dimensional square lattice consisting of $N$ sites. The stress at each site increases uniformly and linearly in time at a constant rate ($v_{0}$). When the stress at a particular site reaches a given threshold value ($S_{c}$), an earthquake event is triggered. We choose $D_{1}$ as a constant stress threshold. Sites with stress greater than $S_{c}-D_{1}$ are identified and grouped into clusters of nearest neighbors. Only the cluster containing the triggering site participates in the earthquake. For every site in this cluster, reduce its stress to a new stress value $S_{c}-D_{s}$, where $D_{s}=D_{2}+D_{3}$. Here, $D_{2}$ is a fixed paramater and $D_{3}$ is a random parameter. In this model ”Cluster” defines a set of connected sites (grids) where stress level exceed the choosen threshold and ”Avalanche” defines as a specific cluster which contains the site with maximum stress (epicenter) at that time. If $n$ is the number of sites in this cluster then, the resulting seismic moment ($M$) and earthquake magnitude ($m$) are related as,
$m=\log M=\frac{3}{2}\log n$
@@ II.2 Sandpile-like model of earthquake
which means that, earthquakes involve a stress release, but the variations in stress drop are small relative to the stress level at which events are initiated. We introduce another spatial correlation parameter ($D_{d}$), where the tresholds are correlated within a typical size of $D_{d}^{2}$ (see e.g., [17]). Large $D_{d}$ value will give large avalanche size.
@@ II.3 Inequality indices used
In quantifying the inequality of the earthquake events, in simulations and in the data, we use inequality indices that are traditionally used to quantify inequality of wealth distribution in a population, viz. the Gini [18] and Kolkata [19] indices. In measuring that we need to first construct the Lorenz curve. It is done as follows: For a set of numbers, here representing earthquake energy released, are first sorted in ascending order. Then $L(p)$ is measured, which is the cumulative fraction of the smallest $p$ fraction of the events. Clearly, $L(0)=0$ and $L(1)=1$. This is called the Lorenz curve. If all events were of exactly equal size, then this would be a diagonal line from (0,0) to (1,1). Any departure from the diagonal line is, therefore, a measure of inequality of the events. The Gini index is defined as the area between the diagonal line and the Lorenz curve, normalized by the area under the diagonal line (1/2). Formally, $g=1-2\int\limits_{0}^{1}L(p)dp$. The value of $g$ is always between 0 and 1, representing complete equality and complete inequality, respectively.
The Kolkata index is defined as the largest 1k1-k fraction of the events that accounts for the $k$ fraction of the cumulative energy released in this case. Formally, it can be obtained by solving 1k=L(k)1-k=L(k). It is a generalization of the Pareto’s 80-20 law in wealth (and many other) distributions, that states that 80% of wealth is typically possessed by the 20% of the richest individuals. Here it gets modified as 1k1-k fraction of the largest events accounts for the $k$ fraction of the total energy released.
These measured have been used recently in responses in systems approaching criticality and have been found to have universal features (see e.g., [20]). Here we use these for earthquake statistics (see also [21]).
@@ III Results
It is known that in systems approaching a critical point, the responses tend to approach a diverging point and the corresponding inequality measures $g$ and $k$ becomes nearly equal before the system reaches the critical point [20, 10]. For Self Organized Critical (SOC) systems also, this near universal behavior is seen [9].
@@
Figure 1: The top figure shows the time series plot of the avalanches for the train model. The bottom figure is the same series, but now sorted according to the corresponding Gini index associated with each event (calculated from the immediately preceeding 100 events. The larger avalanches tend to accumate near larger $g$ values.).
Figure 2: The binned avalanches are shown as a function of $g$ values, with a binning range of 0.01 in the $g$ values for the different system sizes in the 1-d train model. A clear postive correlation is observed.
@@
First we apply this to the response statistics of the train model. We choose a window of 100 consecutive events from the time series of the avalanche sizes (mimicking energy release here) and calculate $g$ and $k$. We then assign that $g$ and $k$ values to the event immediately following the 100 event window. We go on to do this assignments as we slide the window in time by one event each time. This means that all events, except the first 100, have an inequality index value assigned to it. Then we take all those labelled events and sort them in ascending order of the $g$ (and $k$) values. Fig. 1 shows the unsorted and the sorted time series. There is a tendency of accumulation of the larger events towards larger $g$ value. A similar trend (not shown) is observed when the sorting is done with respect to the $k$ values.
Figure 3: The top figure shows the time series of the avalanches seen in the 2d SOC model. The bottom figure is the same time series, but now sorted according to the $g$ value associated with each event (calculated from the immediately preceeding 100 events). Similar to the train model, a clear correspondence is noted with high $g$ and large avalanches, except for the vary high $g$ values, where events are less.
Figure 4: The binned avalanches with $g$ values for the 2d SOC model. The bin size is 0.01. A positive correlation with increasing $g$ is observed, as in the case of the train model.
We then also plot, in Fig. 2, the total energy released for a particular $g$ values (binned in an interval of 0.01). It shows a clear increase with $g$, before it suddenly drops for very high $g$ values, where events are very small in number.
Figure 5: Earthquake energy distribution plots of real earthquake data. Earthquake magnitude was converted into energy using the relation $E\propto 10^{1.5M}$, where $M$ denotes the earthquake magnitude. The resulting energy is expressed in joules. This formula is taken from Peter Bormann et al. [23], in this paper they proposed $E=10^{1.5M}+4.8$, but in our method we ignore the constant term as it do not effect the $g$ value, only the magnitude of earthquake energy increases. Exponent value for each region is mention on the right-most figure. The common fitted Power law exponent is $b_{c}=-1.733\pm 0.002$ with KS D (Kolmogorov-Smirnov Distance)=0.166=0.166.
@@
Figure 6: Top row represents earthquake energy in Joule vs Year for different regions, to make these plots we take every 100100 events. Bottom row represents the same events but now sorted according to their $g$ values. The earthquake data are taken from USGS earthquake catalog from year 197520251975-2025. We calculate gkg-k index by using sliding window approach. We convert the ISO time scale into UTC time scale and then apply different window sizes $W=25,50,100,200,400$ and $800$ to calculate gini and Kolkata inequality indices. The calculated gkg-k is assigned to the very next event which is outside the window. These plots show similar trend as those in the model.
Next we do the same exercise for the 2d sandpile like model. In Fig. 3, the unsorted (original time ordered) and the sorted (according to $g$ values) event series are shown. A clear trend of accumulation of larger events towards higher $g$ value is again observed. The plot (Fig. 4) of the energy released within a fixed interval of $g$ values show a peak, typically around 0.87, which is also known to be the near universal trend in SOC models. The curve then drops due to lack of events with very high $g$ values.
@@
Finally, we apply this to the real data of earthquake catalogs in different parts of the world (Southern Japan, South-East Asia, North America and Indonesia). The coordinates used are as follows: Southern part of Japan: Minimum latitude 30.145, Maximum latitude 41.278 Min longitude 129.792, Max longitude 142.119. South East Asia: Min latitude 20.633, Max latitude 35.747 Min longitude 67.588, Max longitude 99 Western part of North America: Min latitude 12.383, Max latitude 61.143 Min longitude (-145.547), Max longitude (-46.318) Indonesia: Min latitude (-7.537), Max latitude 5.703 Min longitude 95.142, Max longitude 107.051. The catalog used (USGS) can be found in [22].
First we check that the data set show Gutenberg-Richter scaling (see Fig. 5). The lower cut-off used in magnitude is 4.5. Then the magnitudes are conveted to energy [23] (inequalities in magnitudes were reported very recently in [21]). In Fig. 6, the unsorted (actual order) and sorted (according to $g$ values) can be seen. Once again, a clear trend of accumulations of larger events towards higher $g$ values are observed. In Fig. 7, the energy released within a fixed range (of interval 0.01) $g$ values are shown. It also shows a clearly increasing trend, as in the case of the models mentioned above.

### v2

@@ Abstract
It was conjectured for a long time that the tectonic plates are in a self-organized state of criticality and that the Gutenberg-Richter law is a manifestation of that. It was recently shown that for a system near criticality, the inequality of their responses due to external driving would sharply rise and show universal behavior that could indicate proximity of the system to a critical point. As a result, measures such as the Gini and Kolkata indices that quantify inequality, can also serve as indicators of imminent criticallity and that of diverging (system spanning) responses. In the context of earthquakes, such a large response would correspond to events of high magnitudes. In this work, we show with numerical simulations and seismic data analysis that large earthquake events have a tendency to follow events that are highly unequal, similar to the case of a system near a critical point. Therefore, a continuous monitoring of the inequality indices of the earthquake time series could be useful for hazard estimates. We have applied this framework to models of earthquakes as well as to the earthquake time series from various tectonically active regions, such as North America, Southern Japan, parts of South-East Asia and Indonesia. The findings also indicate a quantitative estimate of the distance from criticality, when the tectonic plates are viewed as a self-organized critical system.
@@ I Introduction
When tectonic plates move (at a very slow rate), they undergo a stick-slip motion that causes earthquakes when the stored potential energy is released due to a sudden slip event. While much is known through more than a century of studies related to the motions and the subsequent energy released in tectonic plates, its understanding still faces severe difficulties, particularly due to the inaccessibility of the regions of origin (often tens to hundreds of kilometers underground) [1]. However, the phenomenon of earthquake has fascinated researchers from various disciplines alike not only due to their devastating consequences, but also due to their remarkable regularities in the statistical sense [2].
@@
Particularly, it is known that the earthquake magnitudes follow the Gutenberg-Richter law [3], $\log_{10}N=a-bM$, stating that the number of earthquakes ($N$) of magnitude $M$ is related to $M$ exponentially ($a$ and $b$ are constants). It is also observed that the rate of aftershocks decays with time in a power-law $n(t)=K/(t+c)^{p}$, known as the Omori-Utsu law [4, 5]. There are other empirical laws about earthquake dynamics as well, which are verified using data analysis across various regions in the world. While a complete understanding of the mechanism behind the earthquake phenomenon is still lacking, the empirical observations have been useful in gaining insights into the earthquake dynamics, improving the prospective to model the phenomenon and to also aim for a forecasting possibility, which has so far largely evaded complete understanding.
@@
There have been some efforts in searching for regularities in earthquake statistics that could potentially lead to hazard analysis. One such observation is the change in the exponent value ($b$) prior to a large earthquake event [6, 7, 8, 9]. This method follows from the stress dependence of avalanche size exponent in sheared rocks, and also in fracture. In such cases, the avalanche size exponent value decreases prior to a large event. In earthquake also, when expressed in terms of energy, the size distributions follow a power law. The values of the power law exponent (linearly related to the above mentioned $b$-value of GR law) decreases, if the events are sampled prior to a large event (see e.g., [8]). Also, other complexity measures prior to large earthquakes have been studied in recent years with the aim of possible forecasting (see e.g., [12, 10, 11]).
@@
In this work, we analyze a regularity, which depends on the inequality of the earthquake energy releases. Particularly, we note that prior to a large earthquake, the energy released are more unequal than when such large events are not happening. Inequality in responses of a driven disordered SOC-like system is known to be useful in terms of its universal features [13, 14]. It has also shown promises, both in models [15] and (compression) fracture experiments [16], to have forecasting capabilities of imminent large events. Here, we have measured inequality of energy released within a time-window for earthquake time series in various models and also on the data from several earthquake prone regions. It is seen that in most cases, large events follow a sequence of highly unequal events quantified through elevated values of the Gini and Kolkata indices measured on such sequences.
@@
For the remaining part of the paper, we first describe the models used and the method of measuring the inequalities in their responses. We then correlate the inequalities of the responses of those models with subsequent event sizes. Finally, we analyze data from different earthquake prone regions and show that qualitatively such correlations hold there as well, indicating an underlying principle of relating response inequalities with proximity to critical point and/or lowered $b$-value.
@@ II Models and methods
Here, we first describe the models used for this study and then move on to define the inequality indices that are used.
II.1 Sandpile-like model of earthquake
There have been many attempts to model earthquake phenomena throguh cellular automata type models, for example the Olami-Feder-Christensen model [17] etc. In this work, we adopt the model introduced in ref. [18]. Sandpile models typically measure event size by the number of topplings during an avalanche. The present model defines physically meaningful earthquake observables. As a result, a Gutenberg–Richter (Power) law emerges only in a specific parameter regime. We consider a two-dimensional square lattice consisting of $N$ sites. The stress at each site increases uniformly and linearly in time at a constant rate ($v_{0}$). When the stress at a particular site reaches a given threshold value ($S_{c}$), an earthquake event is triggered. We choose $D_{1}$ as a constant stress threshold. Sites with stress greater than $S_{c}-D_{1}$ are identified and grouped into clusters of nearest neighbors. Only the cluster containing the triggering site participates in the earthquake. For every site in this cluster, the stress is reduced reduced to $S_{c}-D_{s}$, where $D_{s}=D_{2}+D_{3}$. Here, $D_{2}$ is a fixed paramater and $D_{3}$ is a random parameter. In this model ”Cluster” defines a set of connected sites (grids) where stress level exceed the choosen threshold and ”Avalanche” defines as a specific cluster which contains the site with maximum stress (epicenter) at that time. If $n$ is the number of sites in this cluster then, it is related to the resulting earthquake magnitude ($M$) as,
$M=\frac{3}{2}\log n.$ (1)
@@ II.1 Sandpile-like model of earthquake
which means that, earthquakes involve a stress release, but the variations in stress drop are small relative to the stress level at which events are initiated. We use another spatial correlation parameter ($D_{d}$), where the tresholds are correlated within a typical size of $D_{d}^{2}$ (see e.g., [19]). A large value of $D_{d}$ will result in a larger spatial correlation in thresholds, consequently resulting in a larger avalanche, on average. The GR and Omori-Utsu laws observed in this model were reported in ref.[18]. Here we look at the inequalities of the avalanche sizes and attempt to extract signals of imminent large events from the behavior of the inequality indices.
II.2 Train model of earthquake
The train model of earthquake is a simplification of the Burridge-Knopoff model [20], where the stick-slip motion of the overlapping tectonic plates are modeled through massive blocks, connected through linear elastic springs, and driven over a rough surface. The frictional force creates the stick-slip motion and the (elastic) energy released can be measured through the motion of the blocks.
We consider a further simplification on the model [21, 22], where we take the frictional surface to be discrete set of blocks, but fixed in space. When the “train” moves over it, slowly pulled by the rightmost block (“engine”), say, some blocks of the train overlap with the fixed blocks below. At that point, they have a random pinning force that oppose the motion, and is drawn from a uniform distribution in (0,1). The forces on a block in the train at any given instance are, therefore, the elastic forces from its nearest neighbors due to stretching of the elastic springs, and the random pinning force. If the pull due to the elastic force exceeds the random pinning, then the block is moved towards the direction of the pulling, hopping over to the next block below, only if that space is not occupied already. The train is restricted to move along the direction of pulling and never opposite.
Figure 1: The top row figures show the time series of the avalanches seen in the 2d SOC model. The bottom row figures are the same time series, but now sorted according to the $g$ value associated with each event that are calculated from the immediately preceeding 100 events. A clear correspondence is noted with high $g$ values and large avalanches, except for the vary high $g$ values, where event numbers are less.
@@ II.3 Inequality indices used
In quantifying the inequality of the earthquake events, in simulations and in the data, we use inequality indices that are traditionally used to quantify inequality of wealth distribution in a population, viz. the Gini [24] and Kolkata [25] indices. In measuring such indices, we need to first construct the Lorenz curve. It is done as follows: A set of numbers, here representing earthquake energy released, are first sorted in the ascending order. Then the Lorenz curve, $L(p)$ is defined as the cumulative fraction of energy released by the smallest $p$ fraction of the events in that sequence. Clearly, $L(0)=0$ and $L(1)=1$. If all events were of exactly equal size, then the Lorenz curve would be a diagonal line from (0,0) to (1,1). Any departure from the diagonal line is, therefore, a measure of inequality of the events. The Gini index is defined as the area between the diagonal line and the Lorenz curve, normalized by the area under the diagonal line (1/2). Formally, $g=1-2\int\limits_{0}^{1}L(p)dp$. The value of $g$ is always between 0 and 1, representing complete equality and complete inequality, respectively.
The Kolkata index is defined as the largest 1k1-k fraction of the events that accounts for the $k$ fraction of the total energy released in the sequence. Formally, it can be obtained by solving 1k=L(k)1-k=L(k). It is a generalization of the Pareto’s 80-20 law in wealth (and many other) distributions, that states that 80% of wealth is typically possessed by the 20% of the richest individuals. Here it gets modified as 1k1-k fraction of the largest events accounts for the $k$ fraction of the total energy released.
These measured have been used recently for systems approaching criticality and have been found to have universal features (see e.g., [26]), particularly near a critical point. Here we use these for earthquake statistics (see also [27]). The codes used for calculating the inequality indices using the different approaches are publicly available in ref. [28].
@@ III Results
It is known that in systems approaching a critical point, the responses tend to diverge (as does the correlation length) and the corresponding inequality measures of those responses, e.g., $g$ and $k$, become nearly equal ($\sim 0.87$) shortly before the systems reach the critical points [26, 14]. For Self Organized Critical (SOC) systems as well, this behavior is seen [13], which is almost independent of the universality class of the model. Recalling that the minimum values for $g$ and $k$ are 0 and $0.5$ respectively, the elevated inequality level of the responses from a system would indicate the proximity of a critical point and hence an imminent large response/event (see e.g., [16] for application to fracture experiments).
@@
Figure 2: The top row figures show the time series plot of the avalanches for the train model. The bottom row figures are the same series, but now sorted according to the corresponding Gini index associated with each event (calculated from the immediately preceeding 100 events). The larger avalanches tend to accumate near larger $g$ values.
Here we first apply this to the 2d sandpile like model. We choose a temporal window size of $W=100$ consecutive events and calculate $g$ and $k$. Then we assign these $g,k$ values to the next event that immediately followed (not included) the events in the window. We then slide the window by one event and repeat the same. In this way, all events in the time series of avalanches of the model (except the first $W$ events) will have a $g,k$ values assigned to it that was calculated from the preceeding $W$ events. If an elevated inequality is observed prior to a large event, the $g,k$ values attached to the larger events should be higher. In Fig. 1, the original time ordered avalanche series of the model and the same series sorted according to $g$ values associated with the events are shown. A clear trend of accumulation of larger events towards higher $g$ value ($\sim 0.87$) is observed. A similar behavior is seen for the sorting according to the $k$ values (see Fig.7). There are other ways to visualize this trend of unequal events preceeding a large one, for example through the time series of the events superposed with the sliding window calculations of $g,k$ (see Fig.8) and some similar other ways (see Figs.9 and 10 and the related discussions there).
@@
Then we apply the same method to the response statistics of the train model. We choose a window of $W=100$ consecutive events from the time series of the avalanche sizes (mimicking energy release here) and calculate $g$ and $k$. We then assign that $g$ and $k$ values to the event immediately following the 100 event window. We go on to do this assignments as we slide the window in time by one event each time. This means that all events, except the first 100, have inequality indices values assigned to it. Then we take all those labeled events and sort them in ascending order of the $g$ (and $k$) values. Fig. 2 shows the unsorted and the sorted time series. There is a tendency of accumulation of the larger events towards larger $g$ value. A similar trend (see Fig.11) is observed when the sorting is done with respect to the $k$ values. As before, the time series of $g,k$ can also be a way to visualize this trend (see Fig.12). The cumulative energy released (see Fig.13) shows the similar trend as in the case of 2d SOC model.
@@
Figure 3: Top row figures represent earthquake energy in Joule vs time for different regions. Bottom row figures represent the same events but now sorted according to their $g$ values. The earthquake data are taken from the USGS earthquake catalog from January 19751975 to October 20252025. We calculate $g,k$ indices by using a sliding window approach (see text). We convert the ISO time scale into UTC time scale and take $W=100$ immediate preceeding an event to calculate the Gini and Kolkata indices for that event. These plots show a similar trend of accumulation of larger events for higher $g$ values as seen in the models.
@@
Finally, we do the same exercise for the real data of earthquake catalogs in different parts of the world (Southern Japan, South-East Asia, North America and Indonesia). The coordinates used are as follows: Southern part of Japan: Minimum latitude 30.145, Maximum latitude 41.278 Min longitude 129.792, Max longitude 142.119. South East Asia: Min latitude 20.633, Max latitude 35.747 Min longitude 67.588, Max longitude 99 Western part of North America: Min latitude 12.383, Max latitude 61.143 Min longitude (-145.547), Max longitude (-46.318) Indonesia: Min latitude (-7.537), Max latitude 5.703 Min longitude 95.142, Max longitude 107.051. The duration is from January 1975 to October 2025. The catalog used (USGS) can be found in [29].
First we check that the data set show Gutenberg-Richter scaling (see Fig.14). The lower cut-off used in magnitude is 4.5. Then the magnitudes are conveted to energy [30] using
$E=10^{1.5M+C},$ (2)
where we have ignored the prefactor $C$ as it does not influence the inequality measures. Note that inequalities in the magnitudes were reported very recently in [27]. In Fig. 3, the unsorted (actual order) and sorted (according to $g$ values) series are shown for earthquake sequences in the different regions. Once again, a trend of accumulations of larger events towards higher $g$ values are observed. A similar trend is observed (see Fig.15) when the events are sorted according to their $k$ values, instead of $g$. The time series plots for $g,k$ along with the events and the cumulative energy plots (see Figs.16 and 17) reflect the similar trends.
Figure 4: The top row figures show the earthquake events in their original order of occurence, the middle and bottom row figures show the same events when arranged in terms of their associated $g$ and $k$ values respectively, where the $g$ and $k$ values for any event is calculated from a dynamic time window defined as all events between the previous large ($M\geq 7.5$) event and the current event (excluding both the ends). A clear correlation between higher inequality and larger magnitude is seen for each of the tectonically active region.
Other than the fixed sized sliding window approach, we also look at a dynamic reset approach as follows: In the earthquake energy data, we continue to measure $g,k$ with a window size increasing with time (i.e., keep the left end fixed and move the right end of the window one event at a time). When a large event (magnitude $\geq 7.5$) is encountered by the right end of the window, the left end is reset to that event and the right end is again moved one step at a time and so on. The method seeks to capture precursory events prior to a large one, which is not necessarily confined within the fixed window size considered before. Once again, higher $g,k$ values are observed prior to large events. Note that the assignment of $g,k$ values to an event is done similarly as before, i.e. it is assigned to the event just following (but outside) the window considered for calculating $g,k$. In Fig. 4, we show the events in their original (temporal) order and then in the increasing order of their $g$ and $k$ values. It shows a remarkable correlation between higher inequality and higher magnitude of earthquakes. In Fig. 5, we plot $k$ vs $g$ that gives a characteristic shape $g=\frac{\ln(1-k)-\ln(k)}{\ln(1-k)+\ln(k)}$ known for a broad class of Lorenz curves $L(p)=1-(1-p)^{\theta}$ (or $L(p)=p^{\theta}$) drawn from any power law event size distribution (and is independent of the power law exponent; see [26]). We have also shown the $g=k$ line to highlight that most of the large events happen around (generally above) it.
Figure 5: The $g$ and $k$ values obtained from the dynamic reset method (see text) are plotted and the largest 20 events are highlighted. All of the largest 20 events happen around $g\approx k$ and for higher values of the inequality indices. The time variations of $g,k$ and the corresponding event series are shown in the Fig.18.
It is important to explore the relation between the elevated values of the inequality indices and a decrease in the so called $b$-value and the prximity to a critical point. It is known that for a power-law size distributed events, say following $P(S)\sim S^{-\beta}$, if the events are arranged in ascending order of their sizes (with $r$ denoting the index of ordering), it will show a diverging function such as $S_{r}\sim(r-r_{0})^{-n}$, where the diverging/system spanning event happens at $r=r_{0}$, with $n=\frac{1}{\beta-1}$ (see e.g., [31]). For a system driven towards a failure/transition, the natural order is often close to the ascending order. For systems with near stationary sequences, precursory events show a growing rate of energy release prior to large events (see e.g., [16]). Therefore, a reasonable assumption is the (nearly) ascending order of events prior to a large event (see e.g., [32]). It is then possible to calculate the Lorenz function and consequently the Gini index in the closed form [14]:
$g(x,n)=1-\frac{2}{1-\left(1-x\right)^{1-n}}\left[1+\frac{\left(1-x\right)^{2-n}-1}{\left(2-n\right)x}\right],$ (3)
where $x=1$ at the critical point. It can be interpreted as a dimensionless parameter indicating proximity to critical point (for example, $T/T_{c}$ in magnetic systems, $\sigma/\sigma_{c}$ in fracture models and so on). A similar interpretation applied for SOC systems even though such parameters are not explicitly tuned in those systems (e.g., some form of normalized height $h/h_{c}$ in sandpiles). In Fig. 6, we plot the variation of $g$ as a function of $n$ (hence, $\beta$) for different values of $x$. Within the plot, we indicate the $g$ values obtained through the reset method mentioned above. Note that an increase in the $g$ value would correspond to either an increase in $n$ (implying a lowering of $\beta$ or $b$-value) or a closer proximity to the critical point (increase in $x$), or both. Therefore, the lowering of $b$-value, observed for many large earthquakes [8], is consistent with increase in the inequality of event sizes prior to a large event. Even though a range in the $n$ values is indicated from the error margin of fitting the GR law from the data, the changes in the exponent could be more than the range, but is generally difficult to estimate for few events (for example, $W=100$ here).
Figure 6: The variations of the Gini index ($g$) are shown with $n$ (hence, the event size distribution exponent $\beta$) for various values of the parameter $x$ that denotes proximity to a critical point ($x=1$ is the critical state; see text for details). The shaded horizontal rectangle denotes the range of $g$ values observed in the dynamic reset (see Fig. 5) method: $\langle g\rangle\pm\sigma_{g}$, where $\langle g\rangle=0.88$ is the average and $\sigma_{g}=0.09$ is the standard deviation. The shaded vertical rectangle denotes the range of $n$ (calculated from the range of $\beta$ obtained in the GR law fit; see Fig.14). The overlapping shaded area is an indicator where the system might residing be during the course of the dynamics, and it is always very close to the critical point, as is conjectured in an SOC picture of tectonic plates.
The intersection area of the two rectangular regions in Fig. 6 indicating the observed $g$ (within one standard deviation) and $n$ (within fitting error bars) values depict the area where the tectonic plate system is likely to reside over the course of the dynamics. The time series plots of $g$ and $k$ evolution and event sizes with time is included in Fig.18. As expected in the case of an attractive fixed point at criticality, a characteristic of SOC dynamics, the system is always very close to the critical point $x=1$.